OpenStudy (lakers):

Fill in the missing term: (x - 8)2 = x2 + ____ + 64. A. 4x B. 8x C. -16x D. 32x

OpenStudy (anonymous):

is that x2 x^2?

OpenStudy (lakers):

yes it is

OpenStudy (anonymous):

(x-8)*2 = x^2 +____ +64 2x-16 = x^2 + ___ + 64 0 = x^2+ ____ - 2x + 64 +16 0 = x^2 +____ -2x + 80

OpenStudy (lakers):

i dont get it?

OpenStudy (anonymous):

i don't either. the left side only has x to the 1st exponent. the right side has an x^2 value. I can't find a way where both sides are equal.

OpenStudy (lakers):

ok thanks